# Analyzing Similarity Metrics for Data Selection for Language Model Pretraining

### v2

@@ 4.2 Pretraining Loss Correlation Results
To measure the ability of embedding models to reflect pretraining loss generalization and to cluster together datapoints with similar difficulty, we report their variance reduction as we vary (1) the number of gradient updates in language model pretraining and (2) as we vary the average cluster size in our K-Means clustering.
We observe that embeddings produced via a small UL2 model (i.e., LM Token Embeds and LM Output Embeds) trained on the same data achieve a much larger reduction in loss variance when compared to off-the-shelf models; these embeddings produce better clusters with datapoints with similar intermediate pretraining losses (Figure 2). These trends hold across multiple different intermediate training checkpoints and for all different cluster sizes. This suggests that training models specifically on the data to perform data curation is preferable to using off-the-shelf embedding models for predicting pretraining loss generalization. Furthermore, we also generally observe that newer embedding models (e.g., Gecko) do not improve upon older alternatives (e.g., BERT and USE), supporting that embedding model improvements in other settings such as retrieval do not translate to this application in the pretraining setting.
@@
To better interpret what is computed by these similarity metrics, we also visualize pairs of datapoints from the various clusters in Figure 1 to understand what types of examples are placed closely together and which have similar pretraining losses. We randomly sample 3 clusters produced by K-Means clusters with the smallest average cluster size with the LM Output Embeds method to extract embeddings. We randomly sample 2 points from these 3 clusters and observe that samples from the same cluster are thematically similar, although they contains very different information (e.g., the last example referring to two different people in history with some governmental relations, although from different times and different countries). We observe that pretraining losses are similar between within-cluster examples, and lower distances correspond to lower differences in pretraining loss.
@@ 4.3 Cluster Purity Results
Figure 3: Comparison of the purity with respect to data source of K-Means clustering produced by various embedding models on the Pile. Cluster purities are averaged over 50 million clusters.
@@
To better understand and interpret the similarity metrics defined by these embedding models, we measure the purity of clusters with respect to the underlying data source. We again remark that none of these embedding models have been trained with knowledge of the data source. Overall, we observe that most embedding models produce fairly pure clusters, where a majority of points come from the same underlying data source (Figure 3). This supports that these embeddings models are generally aligned with human judgement about differences between types of data.
@@
We observe that the embeddings extracted from the small version of the trained downstream language model achieve the highest cluster purity. We also note that Gecko embeddings achieve the third-highest cluster purity, whereas it performs poorly on the task of predicting loss generalization through variance reduction. This suggests that an improved ability to predict pretraining loss generalization cannot simply be explained via producing clusters that are more pure with respect to underlying data sources. Future work could explore incorporating domain information or metadata into the embedding model’s learning objective, as a form of weak supervision (Sam & Kolter, 2023).
Task Baseline USE Gecko BERT LM Token Embeds LM Output Embeds
ARC - Challenge 32.44 32.99 33.73 32.71 32.54 33.53
ARC - Easy 63.78 65.12 64.24 65.08 64.91 65.46
BoolQ 56.47 59.18 58.31 60.08 62.86 61.93
SuperGLUE - CB 42.41 43.45 48.21 41.67 42.86 48.81
SuperGLUE - Copa 75.25 76.33 77.33 74.00 74.67 78.00
HellaSwag 54.98 56.71 57.11 56.54 56.76 57.53
Multi RC 57.71 55.69 55.91 56.27 55.95 55.19
OpenBook QA 46.30 46.20 46.27 46.60 45.93 46.67
PiQA 72.27 72.98 72.74 73.58 72.76 73.63
Race H 38.02 37.94 38.47 38.75 38.84 38.70
Race M 51.71 52.00 53.32 52.88 52.09 52.60
ReCoRD 84.95 84.35 84.64 84.71 84.51 84.91
RTE 54.51 54.87 51.50 54.87 51.62 52.83
Story Cloze 73.93 74.10 74.26 73.95 73.26 74.49
WiC 48.20 47.86 47.49 47.54 48.17 47.28
Winograd 73.99 75.95 77.05 77.29 77.05 76.56
WinoGrande 59.23 58.91 59.01 59.46 59.96 59.25
WSC 74.12 73.92 73.68 74.74 73.22 74.97
Lambada 21.37 29.86 40.12 34.69 31.43 33.82
Natural Questions 10.01 10.76 10.13 10.54 11.13 11.42
SQuAD v2 51.05 54.26 51.70 51.79 58.56 55.37
TriviaQA Wiki 34.39 35.01 32.99 34.95 34.29 36.08
Web Questions 17.05 17.04 17.87 17.98 18.73 19.13
Average 51.92 52.85 53.30 53.07 53.13 53.83
Table 1: Performance of using various embedding models in the diversity-based selection scheme to select a subset of 200B tokens from the Pile to pretrain a 1.7B parameter language model with a UL2 (Tay et al., 2022) objective. Bolding and underlining denote the best and second-best performing methods on each task, respectively. Results are averaged over 3 pretraining runs. We find that embeddings extracted from small version of the downstream language model and Gecko perform the best for pretraining data curation.
4.4 Diversification-based Pretraining Data Curation Results
We report the average downstream task performance of language models trained on pretraining data subsets produced by using various embedding models in the simple diversity-based data curation strategy in Table 1. We first observe that the standardized diversity-based data curation technique applied to all embedding models (both off-the-shelf and those based on small downstream trained language models) outperform the naive random subset selection baseline. Secondly, we observe the LM Output Embeds and Gecko embeddings achieve the best and second-best performance when averaged across all tasks, respectively.
A notable result is that LM Token Embeds performs comparably to Gecko, while outperforming all the other off-the-shelf embedding models. This is of practical interest due to the simplicity of producing these embeddings; they are simply constructed via averaging over the learned token embeddings and, consequently, are extremely quick to compute in comparison to all other embedding methods. Thus, LM Token Embeds serves as a powerful alternative in settings that are compute-limited.
These pretraining results also suggest that even simple notions of similarity, even those that do not account for positional information, are sufficient for many pretraining data curation applications. Finally, they support that the training of models specifically for this task (e.g., on the same pretraining dataset) is preferable to the use of general purpose off-the-shelf embedding models.
@@ 4.4 Diversification-based Pretraining Data Curation Results
Figure 4: Ablation on the number of components in PCA for Gecko and LM Output Embeds. Results are averaged over 50 million sampled clusters from the Pile. Using more components in PCA better cluster points with similar pretraining loss.
@@
Figure 5: Ablation comparing the use of PCA or Random projections for dimensionality reduction in Gecko and LM Output Embeds. Results are averaged over 50 million sampled clusters from the Pile. Dimensionality reduction via PCA performs better for variance reduction when compared to Random Projections.
@@ 4.5 Ablations
We present ablations on the various dimensionality reduction strategies (i.e., technique and resulting size) that we use for the embedding models before running clustering. We present the ablations on Gecko and LM Output Embeds here and defer results on others embeddings to Appendix A.
Impact of Dimensionality on Variance Reduction
The dimensionality of the embeddings used in clustering often must be low for pretraining scales. Here, we run an ablation studying the role of dimensionality (i.e., the number of components in PCA onto which the embeddings are projected) in the variance reduction through clustering (Figure 4). We remark that scaling clustering to accommodate smaller average cluster sizes (i.e., more cluster centers) is intractable for embeddings that have high dimensionality. Thus, we can only report results on various embeddings with a large number (i.e., 256) of components in PCA with larger average cluster sizes. We observe the trend across all embeddings that a higher dimension and larger number of PCA components improves the embedding models’ ability to cluster points by pretraining loss (see other embedding model results in Appendix A.1).
Dimensionality Reduction Techniques
Another common technique to perform dimensionality reduction (especially with respect to maintaining pairwise distances and similarities) is to use random projections. We experiment with using random projections in our loss clustering experiments and observe that it is outperformed by PCA in terms of variance reduction across all cluster sizes (Figure 5).
@@ 5 Discussion
We analyze the fundamental role of computing similarity in pretraining data curation, finding that coarse-grained notions of similarity are particularly useful and that specific models (even simple ones based on averaging learned token embeddings) can outperform other off-the-shelf embedding models. We believe that both our results and the new evaluation framework for similarity metrics provide new insights into the role of similarity in pretraining data curation. For instance, our variance reduction results shows that embedding models tightly cluster points with similar values of pretraining losses, which has larger implications for dynamic data selection. Future work could study that, during training, we can compute the loss on the centroid in each cluster, and use this to select clusters on which to train, that contain points with high values of these intermediate loss. Our method has other, broad potential applications in the pretraining data setting, such as selecting task-specific data for finetuning (Xia et al., 2024) or even in evaluating the generation of synthetic data for (Meng et al., 2022; Sam et al., 2024) by looking at the distance between these potential data to be selected or generated and real data. Finally, our results highlight the benefits for designing models specifically for language model pretraining settings, and our framework provides an evaluation to exactly capture their utility.
@@ B.1 Additional Evaluation Details
We report results averaged over a large numer of downstream tasks. These datasets largely follow two categories: scoring and decoding tasks. Scoring tasks primarly look at the output distribution of the model, while decoding tasks look at text generations from the language model. Scoring tasks are performed as 1-shot (i.e., giving one demonstration of format and answer), while decoding is performed zeroshot. For scoring tasks, we look at the standard top-1 accuracy. The list of scoring tasks is as follows:
  • ARC Challenge: Easy and Challenge (Clark et al., 2018)
  • BoolQ (Clark et al., 2019)
  • SuperGLUE - CB and Copa (Wang et al., 2019)
  • HellaSwag (Zellers et al., 2019)
  • MultiRC (Khashabi et al., 2018)
  • OpenBookQA (Mihaylov et al., 2018)
  • PIQA (Bisk et al., 2020)
  • RACE-H and RACE-M (Lai et al., 2017)
  • ReCoRD (Zhang et al., 2018)
  • RTE (Dagan et al., 2010)
  • Story Cloze (Mostafazadeh et al., 2016)
  • WIC (Pilehvar & Camacho-Collados, 2018)
  • Winograd (Levesque et al., 2012)
  • WinoGrande (Sakaguchi et al., 2021)
  • WSC (Levesque et al., 2012)
@@
  • Lambada (Paperno et al., 2016)
  • Natural Questions (Kwiatkowski et al., 2019)
  • Squad v2 (Rajpurkar et al., 2018)
  • Trivia QA Wiki (Joshi et al., 2017)
  • Web Questions (Berant et al., 2013)
@@ B.3 Additional Hyperparameter Details
For our K-Means clustering, we perform clustering at 4 different levels of granularity in our variance reduction and cluster purity results. We create four sets of clusterings with an average cluster size of 25, 50, 100, 150, with a minimum cluster size of $\frac{1}{5}$ times the average cluster size and a maximum cluster size of $5$ times the average cluster size. For both RAC and K-Means, we use the squared L2 (Euclidean) distance.
@@ B.4 Dimensionality Reduction Details
For our random projections, we use a sparse random projections onto values of $-\frac{\sqrt{n}}{\sqrt{64}},0,\frac{\sqrt{n}}{\sqrt{64}}$ with probabilities $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ respectively (i.e., the default parameters via scikit-learn). We also L2 normalize the result of random projections.

### v3

@@ 4.2 Pretraining Loss Correlation Results
To measure the ability of embedding models to reflect pretraining loss generalization and to cluster together datapoints with similar difficulty, we report their variance reduction as we vary (1) the number of gradient updates performed by the intermediate pretraining checkpoint and (2) as we vary the average cluster size in our K-Means clustering.
We observe that specialized embeddings (i.e., LM Token Embeds and LM Output Embeds) from a small version of the downstream trained model, which is trained on the same pretraining data, achieve a much larger reduction in loss variance when compared to off-the-shelf models. These specialized embeddings produce better clusters with datapoints with similar intermediate pretraining losses, across all cluster sizes (Figure˜3) and for all intermediate pretraining checkpoints (Figure˜3). This suggests that training models specifically on the data to perform data curation is preferable to using off-the-shelf embedding models for predicting pretraining loss generalization. Furthermore, we also generally observe that newer embedding models (e.g., Gecko) do not improve upon older alternatives (e.g., BERT and USE), supporting that embedding model improvements in other settings, such as retrieval do not translate to this application in the pretraining setting.
@@
To better interpret these similarity metrics, we visualize pairs of datapoints from the clusters in Figure˜1 to understand what types of examples are placed closely together and which have similar pretraining losses. We randomly sample 3 clusters produced by K-Means clusters with the smallest average cluster size, with the LM Output Embeds method to extract embeddings. We randomly sample 2 points from these 3 clusters and observe that samples from the same cluster are thematically similar, although containing very different information (e.g., the last pair referring to two different government-related people, although from different times and different countries). We observe that pretraining losses are similar among within-cluster examples, and lower distances correspond to smaller differences in pretraining loss.
4.3 Diversification-based Pretraining Data Curation Results
We report the average downstream task performance of language models trained on pretraining data subsets produced by using various embedding models in the simple diversity-based data curation strategy in Table˜1. We first observe that the standardized diversity-based data curation technique applied to all embedding models (both off-the-shelf and those based on small downstream trained language models) outperforms the naive random subset selection baseline. Secondly, we observe that the LM Output Embeds achieves the best performance when averaged across all tasks.
A notable result is that LM Token Embeds performs comparably to Gecko, while outperforming all other off-the-shelf embeddings. This is of immediate practical interest due to the simplicity of LM Token Embeds; they require no forward pass and, consequently, are extremely quick to compute in comparison to all other considered alternatives. Thus, LM Token Embeds serves as a viable alternative in settings that are compute-limited. This also suggests that even simple notions of similarity, even those that do not account for positional information, are sufficient for many pretraining data curation applications. Finally, they support that models specialized to this task (e.g., trained on the same dataset) often outperform the current, general-purpose, off-the-shelf embedding models. While this does include training an additional model, it is a single-time investment that is amortized across all embedding use cases and is easily offset by the significant improvements in data curation quality.
@@ 4.4 Cluster Purity Results
Figure 4: Comparison of the purity with respect to data source of K-Means clustering produced by various embedding models on the Pile, when averaged over 50 million clusters from the Pile. Specialized embedding models have higher cluster purity scores.
@@
To better understand and interpret the similarity metrics defined by these embedding models, we measure the purity of clusters with respect to the underlying data source. We again remark that none of these embedding models have been trained with knowledge of the data source. Overall, we observe that most embedding models produce fairly pure clusters, where a majority of points come from the same underlying data source (Figure˜4). This supports that these embeddings models are generally aligned with human judgment about differences between types of data.
@@
We observe that the embeddings extracted from the small version of the trained downstream language model achieve the highest cluster purity. We also note that Gecko embeddings achieve the third-highest cluster purity, whereas it performs poorly on the task of predicting loss generalization through variance reduction. This suggests that an improved ability to predict pretraining loss generalization cannot simply be explained via producing clusters that are more pure with respect to underlying data sources. Future work could explore incorporating domain information or metadata into the embedding model’s learning objective, as a form of weak supervision (Sam and Kolter, 2023).
@@
Figure 5: Ablation on the number of components in PCA for Gecko and LM Output Embeds. Results are averaged over 50 million sampled clusters from the Pile. Using more components in PCA better clusters points with similar pretraining loss.
@@
Figure 6: Ablation comparing the use of PCA or Random projections for dimensionality reduction in Gecko and LM Output Embeds. Results are averaged over 50 million sampled clusters from the Pile. Dimensionality reduction via PCA performs better than via Random Projections.
@@ 4.5 Ablations
We present ablations on components of our evaluation framework — specifically in how we have chosen to perform dimensionality reduction (i.e., technique and resulting size), which is used for the embedding models before running clustering. We present the ablations for Gecko and LM Output Embeds and defer results on others embeddings to Appendix A.
Less Components in PCA Does Not Significantly Hurt Performance.
The dimensionality of the embeddings used in clustering often must be low for pretraining scales. Here, we run an ablation studying the role of dimensionality (i.e., the number of components in PCA onto which the embeddings are projected) in the variance reduction through clustering (Figure˜5). We remark that scaling clustering to accommodate smaller average cluster sizes (i.e., more cluster centers) is intractable for embeddings that have high dimensionality. Thus, we can only report results on various embeddings with a large number (i.e., 256) of components in PCA with larger average cluster sizes. We observe the trend across all embeddings that a higher dimension and larger number of PCA components improves the embedding models’ ability to cluster points by pretraining loss (see other embedding model results in Appendix A.1).
Dimensionality Reduction with PCA Outperforms Random Projections.
Another common technique to perform dimensionality reduction (especially with respect to maintaining pairwise distances and similarities) is to use random projections. We experiment with using random projections in our loss clustering experiments and observe that it is outperformed by PCA in terms of variance reduction across all cluster sizes (Figure˜6).
@@ 5 Discussion
We present a new evaluation framework for the understudied area of similarity metrics used in language model pretraining data curation. Using our framework, we show that off-the-shelf embeddings—despite their widespread use—often underperform even simple, specialized alternatives such as an average of token embeddings that requires no forward pass. This also suggests that practitioners should train their own embedding models, and it suffices to train them on a small fraction of the data ($\sim$20%) and at a much smaller parameter scale. While it may seem intuitive that embeddings tailored to the pretraining task would outperform generic ones, this has not been systematically demonstrated and is certainly not yet standard practice. Our framework both surfaces this gap and offers a practical tool for guiding the design of embedding models optimized for data curation, which can lead to significant improvements in data efficiency.
While we believe that our results are fairly general, future work could extend our findings to other pretraining corpora. Beyond evaluation, our findings have broader implications for pretraining workflows—for example, in selecting task-specific finetuning data (Xia et al., 2024), or identifying synthetic examples that resemble real data (Meng et al., 2022; Sam et al., 2024), and even studying the impacts of scaling up embedding models in these tasks. Overall, our results underscore the importance of using task-aligned similarity metrics in pretraining, and our framework provides a foundation for future research in building and optimizing the design of embedding models tailored to this critical step of the language modeling pipeline.
Limitations
One limitation of our work is that the empirical demonstration of our new evaluation framework is restricted to the Pile dataset. We believe this is fairly standard in the language model pretraining data curation literature to focus on a single pretraining corpus (e.g., as is done in Tirumala et al. (2023)), due to the computational costs of both data curation algorithms and pretraining.
@@ B.1 Additional Evaluation Details
We report results averaged over a large number of downstream tasks. These datasets largely follow two categories: scoring and decoding tasks. Scoring tasks primarily look at the output distribution of the model, while decoding tasks look at text generations from the language model. Scoring tasks are performed as 1-shot (i.e., giving one demonstration of format and answer), while decoding is performed zeroshot. For scoring tasks, we look at the standard top-1 accuracy. The list of scoring tasks is as follows:
  • ARC Challenge: Easy and Challenge [Clark et al., 2018]
  • BoolQ [Clark et al., 2019]
  • SuperGLUE - CB and Copa [Wang et al., 2019]
  • HellaSwag [Zellers et al., 2019]
  • MultiRC [Khashabi et al., 2018]
  • OpenBookQA [Mihaylov et al., 2018]
  • PIQA [Bisk et al., 2020]
  • RACE-H and RACE-M [Lai et al., 2017]
  • ReCoRD [Zhang et al., 2018]
  • RTE [Dagan et al., 2010]
  • Story Cloze [Mostafazadeh et al., 2016]
  • WIC [Pilehvar and Camacho-Collados, 2018]
  • Winograd [Levesque et al., 2012]
  • WinoGrande [Sakaguchi et al., 2021]
  • WSC [Levesque et al., 2012]
@@
  • Lambada [Paperno et al., 2016]
  • Natural Questions [Kwiatkowski et al., 2019]
  • Squad v2 [Rajpurkar et al., 2018]
  • Trivia QA Wiki [Joshi et al., 2017]
  • Web Questions [Berant et al., 2013]
@@ B.3 Additional Hyperparameter Details
For our K-Means clustering, we perform clustering at 4 different levels of granularity in our variance reduction and cluster purity results. We create four sets of clusterings with an average cluster size of 25, 50, 100, 150, with a minimum cluster size of $\frac{1}{5}$ times the average cluster size, and a maximum cluster size of $5$ times the average cluster size. For both RAC and K-Means, we use the squared L2 (Euclidean) distance.
@@ B.4 Dimensionality Reduction Details
For our random projections, we use a sparse random projections onto values of $\frac{\sqrt{n}}{\sqrt{64}},0,\frac{\sqrt{n}}{\sqrt{64}}$ with probabilities $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ respectively (i.e., the default parameters via scikit-learn). We also L2 normalize the result of random projections.